# On the Sharp Input-Output Analysis of Nonlinear Systems under Adversarial Attacks

### v1

@@ Appendix B Preliminaries on sub-Gaussian variables
$\mathbb{P}\biggr{(}\sum_{t=0}^{T-1}\|s-\tilde{s}\|_{2}L_{\phi}$ $\sqrt{M}(\tau+1)(\|u_{t}\|_{2}-\mathbb{E}[\|u_{t}\|_{2}])\leq\|s-\tilde{s}\|_{2}L_{\phi}\sqrt{M}(\tau+1)\Theta(T\sigma_{u})\biggr{)}$
$\geq 1-\exp\biggr{(}-\Theta\Bigr{(}\frac{T^{2}\|s-\tilde{s}\|_{2}^{2}L_{\phi}^{2}M(\tau+1)^{2}\sigma_{u}^{2}}{T\|s-\tilde{s}\|_{2}^{2}L_{\phi}^{2}M(\tau+1)^{2}\sigma_{u}^{2}}\Bigr{)}\biggr{)}=1-\exp(-\Theta(T)).$
Note that $\mathbb{E}[\|u_{t}\|_{2}]\leq\Theta(\sigma_{u})$ due to (26). Thus, we can write
$\mathbb{P}\biggr{(}\sum_{t=0}^{T-1}\|s-\tilde{s}\|_{2}L_{\phi}$ $\sqrt{M}(\tau+1)\|u_{t}\|_{2}\leq 2\|s-\tilde{s}\|_{2}L_{\phi}\sqrt{M}(\tau+1)\Theta(T\sigma_{u})\biggr{)}\geq 1-\exp(-\Theta(T)).$ (61)
When $T\geq\Theta(\log(2/\delta))$, the probability in (61) is lower-bounded by $1-\frac{\delta}{2}$. Considering the lower bound of (D) completes the proof. ∎
Now, we achieve that the equation (34) in Theorem D.1 holds for all $s\in\mathbb{S}^{M-1}$ and for all $i\in\{1,\dots,r\}$ within a finite time. To take advantage of Lemma D.6, which states the difference of $\sum_{t}z_{t}^{i}(s)$ depending on $s$, we introduce the important lemma presented in Vershynin [2010].
@@
For $\epsilon>0$, consider a subset $\mathcal{N}_{\epsilon}$ of $\mathbb{S}^{M-1}$, such that for all $s\in\mathbb{S}^{M-1}$, there exists some point $\tilde{s}\in\mathcal{N}_{\epsilon}$ satisfying $\|s-\tilde{s}\|_{2}\leq\epsilon$. The minimal cardinality of such a subset is called the covering number of the sphere and is upper-bounded by $(1+\frac{2}{\epsilon})^{M}$.
The covering number tells that if you select $(1+\frac{2}{\epsilon})^{M}$ number of points which achieve the sufficient positiveness of $\sum_{t}z_{t}^{i}(\cdot)$, and show that the difference of $\sum_{t}z_{t}^{i}(\cdot)$ is small enough within the distance $\epsilon$, then all the points will achieve the positiveness of $\sum_{t}z_{t}^{i}(\cdot)$.
@@
Suppose that Assumptions 2.3 and 2.6 hold, and $\mathbb{P}(\bm{W}_{t}^{(\tau,i)}>0)=\mathbb{P}(\bm{W}_{t}^{(\tau,i)}<0)$ for all $t$ and $i$. Assume that there exists a constant $q>0$ such that $\mathbb{P}(\bm{W}_{t}^{(\tau)}\equiv 0)\geq q$ for all $t$. Consider $\nu:=\frac{\sqrt{M\tau}L_{\phi}\sigma_{u}}{\lambda}$ and $z_{t}^{i}(s)$ in (58). Given $\delta\in(0,1]$, when
$T\geq\Theta\left(\frac{\tau^{2}\nu^{8}}{q^{2}}\left[M\log\left(\frac{\nu}{q}\right)+\log\left(\frac{r}{\delta}\right)\right]\right),$ (62)
@@
$\sum_{t=\tau}^{T-1}z_{t}^{i}(s)$ $\geq\Theta\left(\frac{q\lambda T}{\tau\nu^{3}}\right)>0,\quad\forall s\in\mathbb{S}^{M-1},~{}~{}\forall i\in\{1,\dots,r\}$ (63)
@@
Note that each time instance in $N_{T}$ (see (39)) appears with probability at least $q$. Thus, we have $N_{T}\geq\frac{q(T-\tau)}{2}$ with probability at least $1-\exp(-\Theta(qT))$ due to the Chernoff bound. Naturally, when $T\geq 2\tau$, we have
$\tilde{N}_{T}\geq\Theta\Bigr{(}\frac{qT}{\tau}\Bigr{)}$ (64)
with the probability at least $1-\exp(-\Theta(qT))$. Accordingly, the right-hand side of (59) is lower-bounded by $1-\exp(-\Theta(\frac{qT}{\tau}))-\exp(-\Theta(\frac{q^{2}T}{\tau^{2}\nu^{8}}))$. Thus, it is guaranteed by the union bound and Lemma D.3 that when
$T\geq\Theta\left(\frac{\tau^{2}\nu^{8}}{q^{2}}\log\left(\frac{1}{\delta}\right)\right),$ (65)
the inequality
$\sum_{t=\tau}^{T-1}z_{t}^{i}(s)\geq\frac{1}{2}\Theta\left(\frac{q\lambda T}{\tau\nu^{3}}\right)>0$ (66)
holds for a fixed $s$ with probability at least $1-\frac{\delta}{2}$.
Now, define $\epsilon^{*}=\frac{1}{4}\Theta\bigr{(}\frac{q}{\tau\nu^{4}}\bigr{)}$. From Lemma D.6, for all $s,\tilde{s}\in\mathbb{S}^{M-1}$ satisfying $\|s-\tilde{s}\|\leq\epsilon^{*}$, we have
$\sum_{t=\tau}^{T-1}z_{t}^{i}(s)-\sum_{t=\tau}^{T-1}z_{t}^{i}(\tilde{s})\geq-\Theta(T\epsilon^{*}L_{\phi}\sqrt{M}\tau\sigma_{u})\geq-\frac{1}{4}\Theta\Bigr{(}\frac{qT}{\tau\nu^{3}}\cdot\frac{L_{\phi}\sqrt{M}\tau\sigma_{u}}{\nu}\Bigr{)}=-\frac{1}{4}\Theta\Bigr{(}\frac{q\lambda T}{\tau\nu^{3}}\Bigr{)}.$ (67)
with probability at least $1-\frac{\delta}{2}$, when $T\geq\Theta(\log(1/\delta))$. If we select $\Theta((1+\frac{2}{\epsilon^{*}})^{M})$ points $\{s_{1},\dots,s_{(1+\frac{2}{\epsilon^{*}})^{M}}\}$ satisfying (66) with probability at least $1-\frac{\delta}{2\cdot(1+\frac{2}{\epsilon^{*}})^{M}}$, then it follows that
$\sum_{t=\tau}^{T-1}z_{t}^{i}(s)\geq\frac{1}{2}\Theta\left(\frac{q\lambda T}{\tau\nu^{3}}\right),\quad\forall s\in\hat{S}=\{s_{1},\dots,s_{(1+\frac{2}{\epsilon^{*}})^{M}}\}$ (68)
with probability at least $1-\frac{\delta}{2}$. Then, due to Lemma D.7, every point in $\mathbb{S}^{M-1}$ is within a distance of $\epsilon^{*}$ from at least one point in $\hat{S}$. In turn, by (67), we have
$\sum_{t=\tau}^{T-1}z_{t}^{i}(s)\geq\frac{1}{4}\Theta\left(\frac{q\lambda T}{\tau\nu^{3}}\right)>0,\quad\forall s\in\mathbb{S}^{M-1}$ (69)
holds with probability at least $1-\delta$. Thus, we replace $\delta$ in (65) with $\frac{\delta}{(1+\frac{2}{\epsilon^{*}})^{M}}$ to arrive at
$T\geq\Theta\left(\frac{\tau^{2}\nu^{8}}{q^{2}}\log\left(\frac{(1+\frac{2}{\epsilon^{*}})^{M}}{\delta}\right)\right)$ $=\Theta\left(\frac{\tau^{2}\nu^{8}}{q^{2}}\left[M\log\left(1+\frac{2}{\epsilon^{*}}\right)+\log\left(\frac{1}{\delta}\right)\right]\right)$
$=\Theta\left(\frac{\tau^{2}\nu^{8}}{q^{2}}\left[M\log\left(\frac{\nu}{q}\right)+\log\left(\frac{1}{\delta}\right)\right]\right)$ (70)
Finally, to satisfy (69) for all $i\in\{1,\dots,r\}$ with probability $1-\delta$, it suffices to satisfy (69) with probability $1-\frac{\delta}{r}$ for all $i$. Thus, we substitute $\frac{\delta}{r}$ for $\delta$ in (D) to obtain the conclusion in (62). ∎
@@
Suppose that Assumptions 2.3 and 2.6 hold. Assume that there exists a constant $\tilde{q}>0.5$ such that $\mathbb{P}(\bm{W}_{t}^{(\tau)}\equiv 0)\geq\tilde{q}$ for all $t$. Consider $\nu:=\frac{\sqrt{M\tau}L_{\phi}\sigma_{u}}{\lambda}$ and $z_{t}^{i}(s)$ in (58). Given $\delta\in(0,1]$, when
$T\geq\Theta\left(\frac{\tau^{2}\nu^{8}}{(2\tilde{q}-1)^{2}}\left[M\log\left(\frac{\nu}{2\tilde{q}-1}\right)+\log\left(\frac{r}{\delta}\right)\right]\right),$ (71)
we have
$\sum_{t=\tau}^{T-1}z_{t}^{i}(s)$ $\geq\Theta\left(\frac{(2\tilde{q}-1)\lambda T}{\tau\nu^{3}}\right)>0,\quad\forall s\in\mathbb{S}^{M-1},~{}~{}\forall i\in\{1,\dots,r\}$ (72)
with probability at least $1-\delta$.
Proof.
We first focus on a restricted class of problems where the following distribution of $\bm{W}_{t}^{(\tau)}$, which is constructed via two types of vectors $v_{t}$ and $\tilde{v}_{t}$, is injected into the system:
$\bm{W}_{t}^{(\tau)}\sim\mathbb{I}_{A}\cdot[\mathbb{I}_{B}\cdot v_{t}+(1-\mathbb{I}_{B})\cdot\tilde{v}_{t}],\quad\mathbb{I}_{A}\sim\text{Bernoulli}(2(1-\tilde{q})),$ (73)
$\mathbb{I}_{B}\sim\text{Bernoulli}(0.5),\quad\mathbb{I}_{A}\perp\!\!\!\perp\mathbb{I}_{B},\quad(\mathbb{I}_{A},\mathbb{I}_{B})\perp\!\!\!\perp v_{t},\quad(\mathbb{I}_{A},\mathbb{I}_{B})\perp\!\!\!\perp\tilde{v}_{t}$
[…]
$\mathbb{P}(\bm{W}_{t}^{(\tau)}\equiv 0)\geq 1-2(1-\tilde{q})=2\tilde{q}-1>0,$ (74)
where the last inequality comes from $\tilde{q}>0.5$.
The roles of the two vectors $v_{t}$ and $\tilde{v}_{t}$ are distinct in the sense that the attacker can select any adversarial distribution for $v_{t}$, whereas the distribution of $\tilde{v}_{t}$ is strategically designed in response to $v_{t}$ to guarantee that the aggregate distribution of $\bm{W}_{t}^{(\tau)}$ satisfies the assumption in Theorem D.8: $\mathbb{P}(\bm{W}_{t}^{(\tau,i)}>0)=\mathbb{P}(\bm{W}_{t}^{(\tau,i)}<0)$. This is achievable since the arbitrary distribution $v_{t}$ can be neutralized by $\tilde{v}_{t}$ to satisfy the sign symmetry assumption, given that $\mathbb{I}_{B}$ ensures equal probabilities of occurrence for both vectors. Using this fact together with (74), we can apply Theorem D.8 by replacing $q$ in (62) and (63) with $12\tilde{q}-$ 1 to arrive at (71) and (72). We then design $\tilde{v}_{t}$ in (73) to have an arbitrarily small magnitude at each time $t$, while preserving its direction. The sign symmetry of the distribution remains intact regardless of the magnitude, and thus both (71) and (72) still hold.
Now, consider another distribution
$\bm{W}_{t}^{(\tau)}\sim\mathbb{I}_{A}\cdot\mathbb{I}_{B}\cdot v_{t},$ (75)
which is a consequence of setting the magnitude of $\tilde{v}_{t}$ in (73) to exactly zero. As the magnitude transitions from arbitrarily small to exactly zero, some nonzero instances of $\bm{W}_{t}^{(\tau)}$ become zero. This implies that the lower bound (72) still holds for the distribution (75) since certain values of $z_{t}^{i}(s)$ that were equivalent to $-|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|$ increase to $|s^{T}\Phi(\bm{U}_{t}^{(\tau)})|$ when $\bm{W}_{t}^{(\tau)}=0$ (see (58)). Finally, using the fact that (75) satisfies
$\mathbb{P}(\bm{W}_{t}^{(\tau)}\equiv 0)\geq 1-\mathbb{P}(\mathbb{I}_{A}=1,\mathbb{I}_{B}=1)=1-\mathbb{P}(\mathbb{I}_{A}=1)\mathbb{P}(\mathbb{I}_{B}=1)=1-2(1-\tilde{q})\cdot 0.5=\tilde{q}$
completes the proof. ∎
Corollary D.10.
Suppose that Assumptions 2.3, 2.6, and 2.8 hold. Consider $\nu:=\frac{\sqrt{M\tau}L_{\phi}\sigma_{u}}{\lambda}$ and $z_{t}^{i}(s)$ in (58). Given $\delta\in(0,1]$, when
$T\geq\Theta\left(\frac{\tau^{2}\nu^{8}}{(2(1-p)^{\tau}-1)^{2}}\left[M\log\left(\frac{\nu}{2(1-p)^{\tau}-1}\right)+\log\left(\frac{r}{\delta}\right)\right]\right),$ (76)
we have
$\sum_{t=\tau}^{T-1}z_{t}^{i}(s)$ $\geq\Theta\left(\frac{(2(1-p)^{\tau}-1)\lambda T}{\tau\nu^{3}}\right)>0,\quad\forall s\in\mathbb{S}^{M-1},~{}~{}\forall i\in\{1,\dots,r\}$ (77)
with probability at least $1-\delta$.
Proof.
Under the equation (10) in Assumption 2.8, we have
$\mathbb{P}(\bm{W}_{t}^{(\tau)}\equiv 0)\geq\mathbb{P}(w_{i}\equiv 0,~{}\forall i\in\{t-1,\dots,t-\tau\})=(1-p)^{\tau}.$
Under $p<\frac{1}{2\tau}$, we have $(1-p)^{\tau}>0.5$, which satisfies the assumption of Lemma D.9. Thus, we can replace $\tilde{q}$ in the lemma with $(1-p)^{\tau}$. ∎
In Corollary D.10, we achieve that $\sum_{t}z_{t}^{i}(s)$ is sufficiently positive for all $t$ and $i$ within a time given in (76). Thus, we arrive at the conclusion that when $\bm{x}_{t}^{(\tau)}=0$ for all $t$, $G^{*}$ is the unique solution to the $\ell_{1}$-norm estimator (12) after a finite time due to Lemma D.1.
Before presenting the main theorem, we provide the following useful lemma.
Lemma D.11.
Suppose that Assumptions 2.1, 2.3, and 2.4 hold. Given $\delta\in(0,1],$ when $T\geq\Theta(\log(1/\delta))$,
$\sum_{t=0}^{T-\tau-1}\|x_{t}\|_{2}\leq\Theta\left(\frac{(\sigma_{u}+\sigma_{w})}{1-\rho}\cdot T\right)$ (78)
@@
×$\hskip 85.35826pt\times\exp\Big{(}\theta\Big{[}\|x_{0}\|_{2}-\mathbb{E}[\|x_{0}\|_{2}]+\sum_{i=0}^{T-\tau-3}\|w_{i}\|_{2}-\mathbb{E}[\|w_{i}\|_{2}]\Big{]}\Big{)}\Big{]}$
$\leq\exp(\theta^{2}\Theta(\sigma_{w}^{2}))\cdot\mathbb{E}\Big{[}\exp\Big{(}\theta\Big{[}\|x_{0}\|_{2}-\mathbb{E}[\|x_{0}\|_{2}]+\sum_{i=0}^{T-\tau-3}\|w_{i}\|_{2}-\mathbb{E}[\|w_{i}\|_{2}]\Big{]}\Big{)}\Big{]}$
$\leq\cdots\leq\exp(\theta^{2}\Theta(T\sigma_{w}^{2}))$
for all $\theta\in\mathbb{R}$, where the inequalities come from applying Lemma B.2 to Assumption 2.4. Thus, the sub-Gaussian norm of $\|x_{0}\|_{2}-\mathbb{E}[\|x_{0}\|_{2}]+\sum_{i=0}^{T-\tau-2}\|w_{i}\|_{2}-\mathbb{E}[\|w_{i}\|_{2}]$ is $\Theta(\sqrt{T}\sigma_{w})$. Furthermore, since the sub-Gaussian norm of $\|u_{i}\|_{2}$ is $\sigma_{u}$ due to Assumption 2.3, the sub-Gaussian norm of $\sum_{i=0}^{T-\tau-2}\|u_{i}\|_{2}-\mathbb{E}[\sum_{i=0}^{T-\tau-2}\|u_{i}\|_{2}]$ is $\Theta(\sqrt{T}\sigma_{u})$ by applying Lemmas B.2 and B.3.
Denote the term in (80) as $S_{T}$. Considering the aforementioned sub-Gaussian norms, the sub-Gaussian norm of $S_{T}-\mathbb{E}[S_{T}]$ is $\Theta(\sqrt{T}\frac{\sigma_{w}+\sigma_{u}}{1-\rho})$ due to the triangle inequality and the homogeneity. Due to the property (29a), one arrives at
$\mathbb{P}\left(S_{T}-\mathbb{E}[S_{T}]\leq\Theta\left(\frac{\sigma_{w}+\sigma_{u}}{1-\rho}\cdot T\right)\right)$ $\geq 1-\exp\left(-\Theta\left(\frac{T^{2}(\sigma_{w}+\sigma_{u})^{2}/(1-\rho)^{2}}{T(\sigma_{w}+\sigma_{u})^{2}/(1-\rho)^{2}}\right)\right)$
$=1-\exp(-\Theta(T))$ (81)
@@
$\leq\frac{C}{1-\rho}[\Theta(T\sigma_{w})+\Theta(T\sigma_{u})]=\Theta\left(\frac{\sigma_{w}+\sigma_{u}}{1-\rho}\cdot T\right),$ (82)
where the last inequality is obtained by applying the property (26) to Assumptions 2.3 and 2.4. Combining (81) and (82) yields
$\mathbb{P}\left(S_{T}\leq 2\cdot\Theta\left(\frac{\sigma_{w}+\sigma_{u}}{1-\rho}\cdot T\right)\right)\geq 1-\exp(-\Theta(T))\geq 1-\delta$ (83)
when $T\geq\Theta(\log(1/\delta))$. Recall from (80) that $\sum_{t=0}^{T-\tau-1}\|x_{t}\|_{2}$ is bounded above by $S_{T}$. This completes the proof. ∎
Now, we present our main theorem, which states that the estimation error to identify $G^{*}$ in the system (11) is bounded by $\Theta(\rho^{\tau})$ when using the $\ell_{1}$-norm estimator.
@@
Suppose that Assumptions 2.1, 2.3, 2.4, 2.6, and 2.8 hold. Consider $\nu:=\frac{\sqrt{M\tau}L_{\phi}\sigma_{u}}{\lambda}$. Let $G^{*}$ be the true matrix governing the system (11) and $\hat{G}_{T}$ denote a solution to the $\ell_{1}$-norm estimator given in (12). Given $\delta\in(0,1]$, when
$T\geq\Theta\left(\frac{\tau^{2}\nu^{8}}{(2(1-p)^{\tau}-1)^{2}}\left[M\log\left(\frac{\nu}{2(1-p)^{\tau}-1}\right)+\log\left(\frac{r}{\delta}\right)\right]\right),$ (84)
@@
$\|G^{*}-\hat{G}_{T}\|_{F}\leq\Theta\left(\frac{\rho^{\tau}L\sqrt{r}}{\lambda}\cdot\frac{(\sigma_{u}+\sigma_{w})\tau\nu^{3}}{(1-\rho)(2(1-p)^{\tau}-1)}\right)$ (85)

### v2

@@ Appendix B Preliminaries on sub-Gaussian variables
$\mathbb{P}\biggr(\sum_{t=0}^{T-1}\|Z-\tilde{Z}\|_{F}L_{\phi}$ $\sqrt{M}(\tau+1)(\|u_{t}\|_{2}-\mathbb{E}[\|u_{t}\|_{2}])\leq\|Z-\tilde{Z}\|_{F}L_{\phi}\sqrt{M}(\tau+1)\cdot O(T\sigma_{u})\biggr)$
$\geq 1-\exp\biggr(-\Omega\Bigr(\frac{T^{2}\|Z-\tilde{Z}\|_{F}^{2}L_{\phi}^{2}M(\tau+1)^{2}\sigma_{u}^{2}}{T\|Z-\tilde{Z}\|_{F}^{2}L_{\phi}^{2}M(\tau+1)^{2}\sigma_{u}^{2}}\Bigr)\biggr)=1-\exp(-\Omega(T)).$
Note that $\mathbb{E}[\|u_{t}\|_{2}]=O(\sigma_{u})$ due to (24a). Thus, we can write
$\mathbb{P}\biggr(\sum_{t=0}^{T-1}\|Z-\tilde{Z}\|_{F}L_{\phi}$ $\sqrt{M}(\tau+1)\|u_{t}\|_{2}\leq 2\|Z-\tilde{Z}\|_{F}L_{\phi}\sqrt{M}(\tau+1)\cdot O(T\sigma_{u})\biggr)\geq 1-\exp(-\Omega(T)).$ (54)
When $T=\Omega(\log(2/\delta))$, the probability in (54) is lower-bounded by $1-\frac{\delta}{2}$. Considering the lower bound of (D) completes the proof. ∎
Now, we will achieve that the inequality (31) in Theorem D.1 holds for all $Z\in\mathbb{R}^{r\times M}$ such that $\|Z\|_{F}=1$, after finite time. To take advantage of Lemma D.6, which states the difference of $\sum_{t}\bar{f}_{t}(Z)$ depending on $Z$, we introduce the important lemma presented in Vershynin (2010).
@@
Define $\mathbb{S}^{r\times M-1}:=\{Z\in\mathbb{R}^{r\times M}:\|Z\|_{F}=1\}$. For $\epsilon>0$, consider a subset $\mathcal{N}_{\epsilon}$ of $\mathbb{S}^{r\times M-1}$, such that for all $Z\in\mathbb{S}^{r\times M-1}$, there exists some point $\tilde{Z}\in\mathcal{N}_{\epsilon}$ satisfying $\|Z-\tilde{Z}\|_{2}\leq\epsilon$. The minimal cardinality of such a subset is called the covering number of the sphere and is upper-bounded by $(1+\frac{2}{\epsilon})^{rM}$.
The covering number argument states that if you select $(1+\frac{2}{\epsilon})^{rM}$ number of points which achieve the sufficient positiveness of $\sum_{t}\bar{f}_{t}(Z)$, and show that the difference of $\sum_{t}\bar{f}_{t}(Z)$ is small enough within the distance $\epsilon$, then all the points will achieve the positiveness of $\sum_{t}\bar{f}_{t}(Z)$.
@@
Suppose that Assumptions 2.4 and 2.5 hold. Consider $\nu:=\frac{\sqrt{M\tau}L_{\phi}\sigma_{u}}{\lambda}$ and $\mathbb{S}^{r\times M-1}:=\{Z\in\mathbb{R}^{r\times M}:\|Z\|_{F}=1\}$. Given $\delta\in(0,1]$, when
$T=\Omega\left(\frac{\tau\nu^{8}}{(2(1-p)^{\tau}-1)^{2}}\left[rM\log\left(\frac{\tau\nu}{2(1-p)^{\tau}-1}\right)+\log\left(\frac{1}{\delta}\right)\right]\right),$ (55)
@@
$\sum_{t=\tau}^{T-1}\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}$ $=\Omega\left(\frac{(2(1-p)^{\tau}-1)\cdot\lambda T}{4\nu^{3}}\right)>0,\quad\forall Z\in\mathbb{S}^{r\times M-1}$ (56)
@@
As in the previous lemma, we define $\bar{f}_{t}(Z):=\mathbb{I}_{\pm}\{\bm{W}_{t}^{(\tau)}=0\}\cdot\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}$. Also, define $\epsilon^{*}=\frac{1}{4}O\bigr(\frac{2(1-p)^{\tau}-1}{\tau^{1/2}\nu^{4}}\bigr)$. From Lemma D.6, for all $Z,\tilde{Z}\in\mathbb{S}^{r\times M-1}$ satisfying $\|Z-\tilde{Z}\|_{F}\leq\epsilon^{*}$, we have
$\sum_{t=\tau}^{T-1}\bar{f}_{t}(Z)-\sum_{t=\tau}^{T-1}\bar{f}_{t}(\tilde{Z})\geq-O(T\epsilon^{*}L_{\phi}\sqrt{M}\tau\sigma_{u})$ $\geq-\frac{1}{4}O\Bigr(T\cdot\frac{2(1-p)^{\tau}-1}{\tau^{1/2}\nu^{3}}\frac{L_{\phi}\sqrt{M}\tau\sigma_{u}}{\nu}\Bigr)$
$=-\frac{1}{4}O\Bigr(\frac{(2(1-p)^{\tau}-1)\lambda T}{\nu^{3}}\Bigr).$ (57)
with probability at least $1-\frac{\delta}{2}$, when $T=\Omega(\log(2/\delta))$. If we select $(1+\frac{2}{\epsilon^{*}})^{rM}$ points $\{Z_{1},\dots,Z_{(1+\frac{2}{\epsilon^{*}})^{rM}}\}$ satisfying (45) with probability at least $1-\frac{\delta}{2\cdot(1+\frac{2}{\epsilon^{*}})^{rM}}$, then it follows that
$\sum_{t=\tau}^{T-1}\bar{f}_{t}(Z)=\frac{1}{2}\Omega\left(\frac{(2(1-p)^{\tau}-1)\lambda T}{\nu^{3}}\right),\quad\forall Z\in\hat{Z}=\{Z_{1},\dots,Z_{(1+\frac{2}{\epsilon^{*}})^{rM}}\}$ (58)
with probability at least $1-\frac{\delta}{2}$. Then, due to Lemma D.7, every point in $\mathbb{S}^{r\times M-1}$ is within a distance of $\epsilon^{*}$ from at least one point in $\hat{Z}$. In turn, by (D), we have
$\sum_{t=\tau}^{T-1}\bar{f}_{t}(Z)\geq\frac{1}{4}\Omega\left(\frac{(2(1-p)^{\tau}-1)\lambda T}{\nu^{3}}\right)>0,\quad\forall Z\in\mathbb{S}^{r\times M-1}$ (59)
holds with probability at least $1-\delta$. Thus, we replace $\delta$ in (44) with $\frac{\delta}{2\cdot(1+\frac{2}{\epsilon^{*}})^{rM}}$ to arrive at
$T$ $=\Omega\left(\frac{\tau\nu^{8}}{(2(1-p)^{\tau}-1)^{2}}\log\Bigr(\frac{2(1+\frac{2}{\epsilon^{*}})^{rM}}{\delta}\Bigr)\right)$
$=\Omega\left(\frac{\tau\nu^{8}}{(2(1-p)^{\tau}-1)^{2}}\left[rM\log\left(1+\frac{2}{\epsilon^{*}}\right)+\log\left(\frac{1}{\delta}\right)\right]\right)$
$=\Omega\left(\frac{\tau\nu^{8}}{(2(1-p)^{\tau}-1)^{2}}\left[rM\log\left(\frac{\tau\nu}{2(1-p)^{\tau}-1}\right)+\log\left(\frac{1}{\delta}\right)\right]\right),$ (60)
where we leveraged Lemma D.4 for the last equality. Note that $T=\Omega(\log(2/\delta))$ required for (D) is automatically satisfied with the recovery time (D). This completes the proof. ∎
In Theorem D.8, we achieve that $\sum_{t}\bar{f}_{t}(Z)$ is sufficiently positive after the recovery time given in (55). Thus, we arrive at the conclusion that when $\bm{x}_{t}^{(\tau)}=0$ and $\bm{\epsilon}_{t}=0$ for all $t$, $G^{*}$ is the unique solution to the $\ell_{2}$-norm estimator (12) after finite time due to Lemma D.1.
We will now generalize for the case of nonzero $\bm{x}_{t}^{(\tau)}$ and $\bm{\epsilon}_{t}$. Before presenting the main theorem, we provide the following useful lemma.
@@
Suppose that Assumptions 2.1, 2.5, and 2.6 hold. Given $\delta\in(0,1],$ when $T=\Omega(\log(1/\delta))$,
$\sum_{t=0}^{T-\tau-1}\|x_{t}\|_{2}=O\left(\frac{(\sigma_{u}+\sigma_{w})}{1-\rho}\cdot T\right)$ (61)
@@
$\hskip 85.35826pt\times\exp\Big(\theta\Big[\|x_{0}\|_{2}-\mathbb{E}[\|x_{0}\|_{2}]+\sum_{i=0}^{T-\tau-3}\|w_{i}\|_{2}-\mathbb{E}[\|w_{i}\|_{2}]\Big]\Big)\Big]$
$\leq\exp(\theta^{2}\cdot O(\sigma_{w}^{2}))\cdot\mathbb{E}\Big[\exp\Big(\theta\Big[\|x_{0}\|_{2}-\mathbb{E}[\|x_{0}\|_{2}]+\sum_{i=0}^{T-\tau-3}\|w_{i}\|_{2}-\mathbb{E}[\|w_{i}\|_{2}]\Big]\Big)\Big]$
$\leq\cdots\leq\exp(\theta^{2}\cdot O(T\sigma_{w}^{2}))$
for all $\theta\in\mathbb{R}$, where the inequalities come from applying Lemma B.2 to Assumption 2.6. Thus, the sub-Gaussian norm of $\|x_{0}\|_{2}-\mathbb{E}[\|x_{0}\|_{2}]+\sum_{i=0}^{T-\tau-2}\|w_{i}\|_{2}-\mathbb{E}[\|w_{i}\|_{2}]$ is $O(\sqrt{T}\sigma_{w})$. Furthermore, since the sub-Gaussian norm of $\|u_{i}\|_{2}$ is $\sigma_{u}$ due to Assumption 2.5, the sub-Gaussian norm of $\sum_{i=0}^{T-\tau-2}\|u_{i}\|_{2}-\mathbb{E}[\sum_{i=0}^{T-\tau-2}\|u_{i}\|_{2}]$ is $O(\sqrt{T}\sigma_{u})$ by applying Lemmas B.2 and B.3.
Denote the term in (63) as $S_{T}$. Considering the aforementioned sub-Gaussian norms, the sub-Gaussian norm of $S_{T}-\mathbb{E}[S_{T}]$ is $O(\sqrt{T}\frac{\sigma_{w}+\sigma_{u}}{1-\rho})$ due to the triangle inequality and the homogeneity. Due to the property (25a), one arrives at
$\mathbb{P}\left(S_{T}-\mathbb{E}[S_{T}]\leq O\left(\frac{\sigma_{w}+\sigma_{u}}{1-\rho}\cdot T\right)\right)$ $\geq 1-\exp\left(-\Omega\left(\frac{T^{2}(\sigma_{w}+\sigma_{u})^{2}/(1-\rho)^{2}}{T(\sigma_{w}+\sigma_{u})^{2}/(1-\rho)^{2}}\right)\right)$
$=1-\exp(-\Omega(T))$ (64)
@@
$\leq\frac{C}{1-\rho}[O(T\sigma_{w}+T\sigma_{u})]=O\left(\frac{\sigma_{w}+\sigma_{u}}{1-\rho}\cdot T\right),$ (65)
where the last inequality is obtained by applying the property (24a) to Assumptions 2.5 and 2.6. Combining (64) and (65) yields
$\mathbb{P}\left(S_{T}\leq 2\cdot O\left(\frac{\sigma_{w}+\sigma_{u}}{1-\rho}\cdot T\right)\right)\geq 1-\exp(-\Omega(T))\geq 1-\delta$ (66)
when $T=\Omega(\log(1/\delta))$. Recall from (63) that $\sum_{t=0}^{T-\tau-1}\|x_{t}\|_{2}$ is bounded above by $S_{T}$. This completes the proof. ∎
Now, we present our main theorem, which states that the estimation error to identify $G^{*}$ in the system (8) is bounded by $O(\rho^{\tau})$ when using the $\ell_{2}$-norm estimator.
@@
Suppose that Assumptions 2.1, 2.4, 2.5, 2.6, and 2.8 hold, and that the approximation error vector satisfies $\|\bm{\epsilon}_{t}\|_{2}\leq\bar{\epsilon}$ for all $t$. Consider $\nu:=\frac{\sqrt{M\tau}L_{\phi}\sigma_{u}}{\lambda}$. Let $G^{*}$ be the true matrix governing the system (8) and $\hat{G}_{T}$ denote a solution to the $\ell_{2}$-norm estimator given in (12). Given $\delta\in(0,1]$, when
$T=\Omega\left(\frac{\tau\nu^{8}}{(2(1-p)^{\tau}-1)^{2}}\left[rM\log\left(\frac{\tau\nu}{2(1-p)^{\tau}-1}\right)+\log\left(\frac{1}{\delta}\right)\right]\right),$ (67)
@@
$\|G^{*}-\hat{G}_{T}\|_{F}=O\left(\left(\frac{\rho^{\tau}L}{\lambda}\cdot\frac{\sigma_{u}+\sigma_{w}}{1-\rho}+\frac{\bar{\epsilon}}{\lambda}\right)\cdot\frac{\nu^{3}}{2(1-p)^{\tau}-1}\right)$ (68)